200.0 mL of each of the following are mixed together: 5.0 x 10-3 M NaF, 2.0 x 10-5 M Ca(NO3)2 and 3.0 x 10-3 M Pb(NO3)2. Which of the following is true? Ksp of CaF2 is 3.9 x 10-11 and for PbF2, 3.7 x 10-8.

a. only PbF2 b. both CaF2 and PbF2 c. neither CaF2 nor PbF2 d. only CaF2
Answer #1

Initial volume of each solution = 200.0 mL

Final volume of mixure = 3 * 200.0 = 600.0 mL

NaF (aq. ) -----------> Na+ (aq.) + F- (aq.)

Final [NaF] = [F-] = 5.0 * 10-3 * 200.0 / 600.0 = 1.67 * 10-3 M

Ca(NO3)2 (aq.) -----------> Ca2+ (aq.) + 2 NO3- (aq.)

final [Ca(NO3)2] = [Ca2+] = 2.0 * 10-5 * 200.0 / 600.0 = 6.67 * 10-6 M

Pb(NO3)2 (aq.) ------------> Pb2+ (aq.) + 2 NO3- (aq.)

Final [Pb(NO3)2] = [Pb2+] = 3.0 * 10-3 * 200.0 / 600.0 = 1.0 * 10-3 M

* CaF2 (s) = Ca2+ (aq.) + 2 F- (aq.)

Qsp = [Ca2+][F-]2

Qsp = (6.67 * 10-6)(1.67*10-3)2

Qsp = 1.86 * 10-11 M3

Since Qsp < Ksp (3.9 * 10-11) It does not form precipitate.

**

PbF2 (s) = Pb2+ (aq.) + 2 F- (aq.)

Qsp = [Pb2+][F-]2

Qsp = (1.0*10-3)(1.67*10-3)2

Qsp = 1.67 * 10-9 M3

Since Qsp < Ksp (3.7 * 10-8) , so it does not form precipitate.

(C) Neither CaF2 nor PbF2 form precipitate